# Disentangling Knowledge Representations for Large Language Model Editing

### v1

@@ Appendix F Baselines
Appendix G Implementation Details
@@
We implement our DiKE method with Pytorch. Our experiments are conducted on NVIDIA A800 (80GB) and NVIDIA GeForce RTX 3090 (24GB). Under such hardware configurations, our method took approximately 2 hours, 3.5 hours, and 4.2 hours to train the GPT-2-XL, GPT-J, and LLaMA-3 models, respectively. For knowledge editing tasks, the average processing time per edit was 8 seconds, 21 seconds, and 35 seconds across these three models, respectively.
@@
To train the Knowledge Representation Disentanglement (KRD) module, we construct a dataset comprising 4,722 knowledge triples, covering 1,784 distinct subjects. The dataset is augmented with subject aliases and rewritten prompts. For each training, 20,000 training samples are generated from this dataset. Specifically, we iterate through the dataset and, for each subject $s$, randomly select two sample pairs, $(s,r,o)$ and $(s,r^{\prime},o^{\prime})$, adding them to the training data until 20,000 samples were created. The module is trained for 5 epochs with a learning rate of $5\times 10^{-5}$. The weighting coefficients for the Knowledge Disentangling Loss, Knowledge Constraint Loss, and Knowledge Reconstruction Loss are set to 1, 0.2, and 1, respectively. The temperature parameter for the Knowledge Disentangling Loss is set to 0.1. The batch size is configured as 4 for GPT2-XL and 16 for GPT-J and LLaMA3. Following MEMIT [Meng et al., 2023], the subject representation is extracted from the last token of the subject. For GPT2-XL, GPT-J, and LLaMA3, subject representations are extracted from layers 17, 8, and 8, respectively. Relation representations are obtained from the last token of the prompt, with extraction from layers 37, 18, and 23 for GPT2-XL, GPT-J, and LLaMA3, respectively.
@@
For the Disentanglement-based Knowledge Edit (DKE) module, the editing layers are selected correspond with the layers from which the subject representations are extracted. To optimize the target knowledge-related representations, the AdamW optimizer [Loshchilov and Hutter, 2019] is used with a learning rate of $5\times 10^{-1}$ for GPT2-XL and GPT-J, and $1\times 10^{-2}$ for LLaMA3. To mitigate overfitting, early-stopping is applied when the loss falls below $5\times 10^{-2}$. For other baselines, experiments are conducted using the code provided by MEMIT [Meng et al., 2023], ensuring all settings, including the hyperparameters, remain consistent with those reported in [Meng et al., 2022, 2023]. All reported results are averaged over 5 runs with different random seeds.
Appendix H Additional Experiments
H.1 Performance Comparison on COUNTERFACT Using GPT2-XL
@@
Table 7: Performance comparison DiKE in terms of Efficacy Score (%), Paraphrase Score (%), and Neighborhood Score (%). The Avg. (%) is the harmonic mean of the three evaluation metrics.
@@
H.2 Effect of Fine-Grained Irrelevant Facts in the KRD Module
@@
H.3 Effect of Training Sample Size in the KRD Module
@@
H.4 Visualization Analysis of Disentangled Representations
@@
H.5 Case Study
Figure 9: LLaMa3 (8B) generation examples of DiKE, ROME and MEMIT. Prompts are italic and green parts in the generation outputs are related to the relational locality answers. Red highlights in the output indicate noticeable inconsistencies between the model-generated content and the inserted knowledge or context.
In this section, we present several generation examples on LLaMA3(8B) utilizing three knowledge editing models: DiKE, AlphaEdit and ROME, to demonstrate the efficacy of knowledge editing through representation disentanglement on FINE-KED. These examples illustrate the models’ abilities to preserve fine-grained neighboring irrelevant knowledge. The generation examples are shown in Figure 9.
@@ H.5 Case Study
Example B. In this example, the new knowledge “The place of birth of Christian Atsu is Klara Church Parish” was inserted into the model. In response to the prompt “The name of the country of citizenship of Christian Atsu," DiKE correctly identified “Ghana” as Christian Atsu’s country of citizenship, based on his professional background. However, after edit by ROME, the model failed to recall the original knowledge related to the neighborhood prompt. Similarly, the model edited by AlphaEdit also failed to provide the original response.

### v2

@@ Appendix E Baselines
Appendix F Implementation Details and Scalability Analysis
F.1 Implementation Details
@@
We implement our DiKE method with Pytorch. Our experiments are conducted on NVIDIA A800 (80GB) and NVIDIA GeForce RTX 3090 (24GB). Under such hardware configurations, our method took approximately 2 hours, 3.5 hours, and 4.2 hours to train the GPT-2-XL, GPT-J, and LLaMA-3 models, respectively. For knowledge editing tasks, the average processing time per edit was 2.79 seconds, 8.72 seconds, and 11.34 seconds across these three models, respectively.
@@
To train the Knowledge Representation Disentanglement (KRD) module, we construct a dataset comprising 4,722 knowledge triples, covering 1,784 distinct subjects. The dataset is augmented with subject aliases and rewritten prompts. For each training, 20,000 training samples are generated from this dataset. Specifically, we iterate through the dataset and, for each subject $s$, randomly select two sample pairs, $(s,r,o)$ and $(s,r^{\prime},o^{\prime})$, adding them to the training data until 20,000 samples were created. The module is trained for 5 epochs with a learning rate of $5\times 10^{-5}$. The weighting coefficients for the Knowledge Disentangling Loss, Knowledge Constraint Loss, and Knowledge Reconstruction Loss are set to 1, 0.2, and 1, respectively. The temperature parameter for the Knowledge Disentangling Loss is set to 0.1. The batch size is configured as 4 for GPT2-XL and 16 for GPT-J and LLaMA3. Following MEMIT (Meng et al., 2023), the subject representation is extracted from the last token of the subject. For GPT2-XL, GPT-J, and LLaMA3, subject representations are extracted from layers 17, 8, and 8, respectively. Relation representations are obtained from the last token of the prompt, with extraction from layers 37, 18, and 23 for GPT2-XL, GPT-J, and LLaMA3, respectively.
@@
For the Disentanglement-based Knowledge Edit (DKE) module, the editing layers are selected correspond with the layers from which the subject representations are extracted. To optimize the target knowledge-related representations, the AdamW optimizer (Loshchilov and Hutter, 2019) is used with a learning rate of $5\times 10^{-1}$ for GPT2-XL and GPT-J, and $1\times 10^{-2}$ for LLaMA3. To mitigate overfitting, early-stopping is applied when the loss falls below $5\times 10^{-2}$. For other baselines, experiments are conducted using the code provided by MEMIT (Meng et al., 2023), ensuring all settings, including the hyperparameters, remain consistent with those reported in (Meng et al., 2022; 2023). All reported results are averaged over 5 runs with different random seeds.
Table 8: Average runtime (seconds) comparison under varying batch sizes on LLaMA-3.
Method BS=1 (s) BS=2 (s) BS=4 (s) BS=8 (s)
MEMIT 9.56 16.47 30.25 59.51
AlphaEdit 8.10 15.07 28.88 58.20
DiKE (Ours) 11.34 18.39 32.82 63.24
Evaluation Criteria: 1. Very Similar (score 7-10): If the two relationships describe entities of the same type or are very close, differing only in details (e.g., ”capital” and ”largest city”). 2. Moderately Similar (score 4-6): If the two relationships describe different types of entities but are in the same domain or background, or have some overlap in the entities they describe (e.g., ”mother” and ”place of birth”). 3. Completely Not Similar (score 0-3): If the two relationships describe entirely different types of entities from different domains or categories, with almost no relation (e.g., ”member of sports team” and ”head of state”). Example Inputs and Outputs: 1. Example 1: Relationship 1: r1 = currency (currency used by item) Relationship 2: r2 = capital (seat of government of a country, province, state or other type of administrative territorial entity) Output: Score: 4 Explanation: These two relationships describe different types of entities—one is about currency, and the other is about political/geographic entities. Although both relate to countries, they differ significantly in their descriptions, so they are moderately similar. 2. Example 2: Relationship 1: r1 = place of birth (most specific known birth location of a person, animal or fictional character) Relationship 2: r2 = spouse (the subject has the object as their spouse (husband, wife, partner, etc.)) Output: Score: 2 Explanation: These two relationships describe completely different things—one is about a birthplace, and the other is about a marriage relationship. There is little to no overlap, so they are completely not similar. 3. Example 3: Relationship 1: r1 = head of state (the chief public representative of a country) Relationship 2: r2 = head of government (the person in charge of running the government of a country) Output: Score: 9 Explanation: These two relationships describe very similar entities—both refer to the highest leaders of a country, with ”head of state” focusing on ceremonial roles and ”head of government” focusing on executive responsibilities. They are very similar. Your task: Relationship 1: r1 = {relation_A} Relationship 2: r2 = {relation_B}
F.2 Scalability Analysis
This section analyzes the scalability of DiKE from multiple perspectives, including parameter efficiency, batch processing behavior, editing cost, and empirical runtime performance.
Fixed Parameter Efficiency. The Knowledge Representation Disentanglement (KRD) module is designed to be lightweight, consisting solely of fixed projection matrices ($\mathbf{W}_{1}$ to $\mathbf{W}_{4}$ in Eq. (5)). As described in the methodology (§3.1), the module operates on the subject representation $\mathbf{h}_{s}$ and relation representation $\mathbf{h}_{r}$, which are extracted from the last-token hidden states of the subject $s$ and the prompt $p(s,r)$ at a specific transformer layer. Importantly, the KRD module does not store subject-specific or relation-specific parameters. Its parameter size therefore remains constant regardless of the number of edited facts, ensuring stable memory usage even as the scale of the underlying knowledge grows.
Batch Scalability Inherited from MEMIT. DiKE retains the robust computational structure and batch scalability of MEMIT. By comparing MEMIT’s solution (Eq. (4)) with DiKE’s closed-form solution (Eq. (19)), it is evident that both methods explicitly solve a least-squares problem for the parameter update. The core computational load is therefore identical. The only additional step introduced by DiKE is the calculation of the disentangled target value $\mathbf{v}^{*}$ via the KRD operation. Since this operation involves only a single forward pass through lightweight matrices ($\mathbf{W}_{1}$ to $\mathbf{W}_{4}$), the added time complexity is negligible. Therefore, theoretically, DiKE inherits the same batch processing capabilities and scalability properties as MEMIT.
Amortized Training Cost. The training of the KRD module constitutes a one-time offline pre-training phase. Once trained, the module is frozen and can be directly applied to any subsequent editing process without requiring retraining or fine-tuning for individual edit samples. This decoupling of training and inference ensures that the online editing latency remains minimal.
Empirical Efficiency Analysis. To empirically assess runtime efficiency, we conduct DiKE, MEMIT, and AlphaEdit under varying batch sizes on LLaMA-3 following the configuration used in Figure 4. Table 8 reports the average runtime for each method across different batch sizes. Across all settings, DiKE exhibits only a small and nearly constant overhead (approximately 2–4 seconds) compared to the baselines. These results indicate that the disentanglement computation introduced by KRD does not form a performance bottleneck and that DiKE maintains stable runtime behavior as the batch size increases.
Appendix G Additional Experiments
G.1 Performance Comparison on COUNTERFACT Using GPT2-XL
@@
Table 9: Performance comparison in terms of Efficacy Score (%), Paraphrase Score (%), and Neighborhood Score (%). The Avg. (%) is the harmonic mean of the three evaluation metrics.
@@
G.2 Performance Comparison on MQuAKE-3K
To investigate DiKE’s generalization capability in more complex scenarios, we have conducted evaluations on multi-hop reasoning using MQuAKE benchmarks. As shown in Table 10, DiKE demonstrates strong performance compared to representative methods such as ROME, MEMIT, and AlphaEdit. This competitive performance can be attributed to our disentanglement mechanism, which effectively isolates the injection of new knowledge from the preservation of irrelevant knowledge. These results demonstrate that our DiKE does not sacrifice generalization performance to maintain fine-grained irrelevant knowledge.
Table 10: Performance comparison of multi-hop editing on MQuAKE in terms of Efficacy Score (%).
Method Avg. 2-hops 3-hops 4-hops
Llama3 29.58 19.79 40.73 27.43
ROME 41.28 40.73 47.30 32.16
MEMIT 33.86 26.37 43.25 30.86
AlphaEdit 40.00 35.68 48.50 33.47
DiKE 44.39 41.62 52.88 35.48
G.3 Effect of Fine-Grained Irrelevant Facts in the KRD Module
@@
G.4 Effect of Training Sample Size in the KRD Module
@@
G.5 Visualization Analysis of Disentangled Representations
@@
Table 11: Performance comparison on FINE-KED under different In-Distribution (ID) and Out-of-Distribution (OOD) settings.
Setting Method Effi. R-Loc.(Easy) R-Loc (Mid.) R-Loc (Hard)
Sub-ID & Rel-ID ROME 99.7 47.6 49.9 35.0
MEMIT 99.5 62.4 58.1 65.8
AlphaEdit 98.0 61.3 52.6 61.2
DiKE 99.3 65.2 70.5 70.8
Sub-ID & Rel-OOD ROME 100.0 59.8 52.2 34.1
MEMIT 97.2 66.9 50.1 50.9
AlphaEdit 94.3 75.0 62.2 58.6
DiKE 98.7 77.4 60.0 61.1
Sub-OOD & Rel-ID ROME 99.9 56.7 50.0 50.7
MEMIT 98.8 65.1 53.8 64.1
AlphaEdit 98.7 67.8 58.8 66.8
DiKE 99.0 72.3 64.1 73.1
Sub-OOD & Rel-OOD ROME 99.7 57.1 54.2 46.0
MEMIT 98.6 64.3 60.0 61.5
AlphaEdit 97.8 68.3 61.6 69.3
DiKE 99.2 73.0 70.3 72.2
G.6 Generalization analysis of KRD Module
To provide a more fine-grained and rigorous assessment of the KRD module’s generalization capabilities, we split the evaluation samples of FINE-KED into four groups based on whether their subjects and relations appeared in the KRD pre-training data: (a) Sub-OOD & Rel-OOD: both subject and relation unseen; (b) Sub-OOD & Rel-ID: unseen subjects, seen relation types; (c) Sub-ID & Rel-OOD: seen subjects, unseen relation types; (d) Sub-ID & Rel-ID: both seen during pre-training.
The comparative results are summarized in Table 11. As observed, DiKE consistently outperforms baseline editors across all settings. Notably, in the most challenging Sub-OOD & Rel-OOD subset, DiKE achieves a Relational Locality (Hard) score of 72.2%, significantly surpassing ROME (46.0%) and MEMIT (61.5%). This confirms that the KRD module maintains strong disentanglement capabilities even when encountering entirely new knowledge patterns.
G.7 Case Study
Figure 9: LLaMa3 (8B) generation examples of DiKE, ROME and AlphaEdit. Prompts are italic and green parts in the generation outputs are related to the relational locality answers. Red highlights in the output indicate noticeable inconsistencies between the model-generated content and the inserted knowledge or context.
In this section, we present several generation examples on LLaMA3(8B) utilizing three knowledge editing models: DiKE, ROME and AlphaEdit, to demonstrate the efficacy of knowledge editing through representation disentanglement on FINE-KED. These examples illustrate the models’ abilities to preserve fine-grained neighboring irrelevant knowledge. The generation examples are shown in Figure 9.
@@ G.7 Case Study
Example B. In this example, the new knowledge “The place of birth of Christian Atsu is Klara Church Parish” was inserted into the model. In response to the prompt “The name of the country of citizenship of Christian Atsu,” DiKE correctly identified “Ghana” as Christian Atsu’s country of citizenship, based on his professional background. However, after edit by ROME, the model failed to recall the original knowledge related to the neighborhood prompt. Similarly, the model edited by AlphaEdit also failed to provide the original response.
@@
Furthermore, we select two representative cases, Example A and Example E, from Sub-OOD & Rel-OOD subset (Appendix G.6) to illustrate that DiKE remains stable and avoids corruption even when both components are unseen during pre-training of KRD module.
Example D. In this case, the new knowledge “The place of burial of Eldar Zeynalov is United States Assistant Attorney General” was injected into the model. When prompted with the neighborhood query “The name of the country of citizenship of Eldar Zeynalov,” DiKE successfully preserved the original factual information and generated the correct response “Azerbaijan.” In contrast, ROME severely distorted the surrounding knowledge and incorrectly produced the injected content as the answer, returning “United States Assistant Attorney General.” AlphaEdit also failed to retain the original information and incorrectly generated “United States of America” as the country of citizenship.
Example E. In this example, the new knowledge “The name of the head of government of Oman is Amsterdam” was inserted. When evaluating the neighborhood knowledge “The name of the capital city of Oman is,” DiKE accurately recalled the correct answer “Muscat” and remained unaffected by the newly edited information. However, ROME incorrectly replaced the capital with the injected knowledge and returned “Amsterdam,” demonstrating a failure to disentangle irrelevant facts. AlphaEdit also failed the query, providing an abnormal multiple-choice-style output that repeatedly listed “Amsterdam” as the supposed capital of Oman.